# View-Invariant Pixelwise Anomaly Detection in Multi-object Scenes with Adaptive View Synthesis

### v1

@@ V-B Implementation Details
V-C Ablation Studies
V-C1 OmniAD architecture
OmniAD has two major changes from RD. First we change the backbone to ResNeXt and second we introduced branching student attention modules $A^{i}$. To test if our changes are significant we test OmniAD with and without ResNeXt (OmniAD-R) and without $A^{i}$ (OmniAD-$A^{i}$) against RD. As shown in the $F_{1}$ scores on ToyCity Tables III and MAD Tables V we can see a consistent performance increase when $A^{i}$ is used.
V-C2 NVS Augmentation
We test the efficacy of our NVS augmentation strategies by evaluating our metrics with and without each of the augmentations. Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table II.
V-D Quantitative Evaluation
In our assessment of diverse AD methods on the ToyCity datasets, we observe a consistent pattern of performance increase OmniAD, OmniAD without ResNext (OmniAD w/o R), and RD when a novel view selection strategy is employed for $F_{1}$ scores on ToyCity and MAD. A similar trend can also be seen for AUROC on both ToyCity and MAD Tables IV,VI. Additional experiments on different scenes is provided in the supplemental material.
@@ V-D Quantitative Evaluation
As mentioned metrics such as pixel-level AUROC can often show very high numbers for datasets such as ToyCity and MAD due to the low number of pixels that may be an anomaly. This can be seen for both ToyCity Table IV and MAD Table VI. Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
TABLE II: Four variations are tested for each of the scenes present in ToyCity. Variation v1 is un-augmented, v2 is with the Query-Aligned Views, v3 is with the interpolated Non-Anomalous views, and v4 is the combination of Query-Aligned Baseline views and Interpolated Non-Anomalous views.
@@
V-E Qualitative Evaluation
Figure9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone. A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score. Three primary failure scenarios are evident. First, as depicted in the ToyCity 1 and 4 anomaly instances, RD may erroneously identify arbitrary areas as anomalous. Secondly, RD has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly. Lastly, RD often overestimates the extent of an anomaly, erroneously marking larger areas than necessary. Similar defects in RD can also be seen from the MAD dataset in which it misses an anomaly or overestimates an anomalous region. Additional samples from OmniAD for each dataset is presented in Figure 10, showing a heat map of the anomaly score as well as the thresholded segmentation.
@@ V-E Qualitative Evaluation
TABLE III: Comparison of Pixel $F_{1}$ scores on ToyCity 1, 2, 3, and 4
Method ToyCity 1 ToyCity 2 ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.432 0.450 0.497 0.496 0.348 0.599 0.491 0.577 0.281 0.384 0.319 0.366 0.379 0.541 0.390 0.523
OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
TABLE IV: Comparison of Pixel AUROC scores on ToyCity 1, 2, 3, and 4
Method ToyCity 1 ToyCity 2 ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.990 0.996 0.999 0.999 0.991 0.998 0.999 0.999 0.979 0.993 0.997 0.999 0.982 0.993 0.995 0.998
OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
TABLE V: Comparison of Pixel $F_{1}$ scores on MAD-Real [8] of Puppy, Pig, and Elephant
@@
TABLE VI: Comparison of Pixel AUROC scores on MAD-Real [8] of Puppy, Pig, and Elephant
@@
VI Conclusion
Our study introduced and addressed a specific anomaly detection (AD) problem we term Scene AD. Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in inspections of uncontrolled environments. To challenge and measure the effectiveness of state-of-the-art anomaly detection for Scene AD, we introduced the ToyCity dataset. This new benchmark focuses on viewpoint variance in scenes of multiple objects and serves as a testbed for our new method OmniAD. We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset, establishing a new state-of-the-art for this task. Particularly, we found that the incorporation of a student attention module increases the efficacy of reverse distillation for pixel-level AD by expanding the effective receptive filed (ERF). Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD. We hope that our work will catalyze further advancements, spurring the creation of advanced anomaly detection techniques capable of managing the complexities associated with varying viewpoints in 3D scenes.

### v2

@@ 5.2 Implementation Details
5.3 Ablation Studies
We conducted ablation studies to validate both the OmniAD architecture and the proposed NVS augmentation strategy.
5.3.1 OmniAD architecture
OmniAD has two major changes from RD. First we change the backbone to ResNeXt and second we introduced branching student attention modules $A^{i}$. To test if our changes are significant we test OmniAD with and without ResNeXt (OmniAD-R) and without $A^{i}$ (OmniAD-$A^{i}$) against RD. As shown in the $F_{1}$ scores on ToyCity Tables 3 and 4 and MAD Tables 7 we can see a consistent performance increase when $A^{i}$ is used.
5.3.2 NVS Augmentation
We test the efficacy of our NVS augmentation strategies by evaluating our metrics with and without each of the augmentations. Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table 2.
5.4 Quantitative Evaluation
In our assessment of diverse AD methods on the ToyCity datasets, we observe a consistent pattern of performance increase OmniAD, OmniAD without ResNext (OmniAD w/o R), and RD when a novel view selection strategy is employed for $F_{1}$ scores on ToyCity and MAD. A similar trend can also be seen for AUROC on both ToyCity and MAD Tables 5,6,8. Additional experiments on different scenes is provided in the supplemental material.
@@ 5.4 Quantitative Evaluation
As mentioned metrics such as pixel-level AUROC can often show very high numbers for datasets such as ToyCity and MAD due to the low number of pixels that may be an anomaly. This can be seen for both ToyCity Table 5,6 and MAD Table 8. Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
Table 2: Four variations are tested for each of the scenes present in ToyCity. Variation v1 is un-augmented, v2 is with the Query-Aligned Views, v3 is with the interpolated Non-Anomalous views, and v4 is the combination of Query-Aligned Baseline views and Interpolated Non-Anomalous views.
@@
5.5 Qualitative Evaluation
Figure  9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone. A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score. Three primary failure scenarios are evident. First, as depicted in the ToyCity 1 and 4 anomaly instances, RD may erroneously identify arbitrary areas as anomalous. Secondly, RD has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly. Lastly, RD often overestimates the extent of an anomaly, erroneously marking larger areas than necessary. Similar defects in RD can also be seen from the MAD dataset in which it misses an anomaly or overestimates an anomalous region. Additional samples from OmniAD for each dataset is presented in Figure 10, showing a heat map of the anomaly score as well as the thresholded segmentation.
@@ 5.5 Qualitative Evaluation
Table 3: Comparison of Pixel $F_{1}$ scores on ToyCity 1 and 2
Method ToyCity 1 ToyCity 2
No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.432 0.450 0.497 0.496 0.348 0.599 0.491 0.577
OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503
OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357
RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390
CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082
FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128
CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081
DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039
Table 4: Comparison of Pixel $F_{1}$ scores on ToyCity 3 and 4
Method ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.281 0.384 0.319 0.366 0.379 0.541 0.390 0.523
OmniAD-R 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
OmniAD-$A^{i}$ 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
RD 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
CFLOW 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
FastFlow 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
CFA 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
DRÆM 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
Table 5: Comparison of Pixel AUROC scores on ToyCity 1 and 2
Method ToyCity 1 ToyCity 2
No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.990 0.996 0.999 0.999 0.991 0.998 0.999 0.999
OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999
OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997
RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998
CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868
FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952
CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984
DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725
Table 6: Comparison of Pixel AUROC scores on ToyCity 3 and 4
Method ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.979 0.993 0.997 0.999 0.982 0.993 0.995 0.998
OmniAD-R 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
OmniAD-$A^{i}$ 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
RD 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
CFLOW 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
FastFlow 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
CFA 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
DRÆM 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
Table 7: Comparison of Pixel $F_{1}$ scores on MAD-Real [13] of Puppy, Pig, and Elephant
@@
Table 8: Comparison of Pixel AUROC scores on MAD-Real [13] of Puppy, Pig, and Elephant
@@
6 Limitations and Future Work
OmniAD and the proposed ToyCity dataset address unsupervised Scene Anomaly Detection in multi-view scenarios, yet several limitations remain. First, each ToyCity image contains at most one anomaly at a time, even though multiple anomalies were added and removed from each scene. In contrast, real-world settings such as disaster assessments or industrial inspections may involve multiple anomalies within a single capture. Note that ToyCity does include partial occlusions, and OmniAD itself does not strictly limit detection to a single anomaly region. Second, although ToyCity was captured with real cameras at varying angles and lighting conditions, its tabletop-scale scenes do not fully encompass the complexities of large scale infrastructure.
Furthermore, our data augmentation assumes sufficient coverage for camera localization. In highly variable environments or when dealing with reflective surfaces, generating high-quality synthetic views can be challenging without accurate camera localization. Nonetheless, as shown in our experiments, OmniAD’s performance does not hinge solely on these augmented views.
In future work, we plan to:
  • 1.
    Extend ToyCity to include multiple anomalies in a single image and introduce dynamic elements such as moving vehicles or pedestrians.
  • 2.
    Validate OmniAD on larger real-world datasets (e.g., complete bridge or ship inspections) to further assess its robustness.
  • 3.
    Investigate faster 3D reconstruction, localization, and noval view synthesis methods.
  • 4.
    Identifying the underlying failure modes in GAN/autoencoder based AD methods under significant viewpoint variation remains an open research question, which we leave for future work.
7 Conclusion
Our study introduced and addressed a specific anomaly detection (AD) problem we term Scene AD. Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in applications in the built environment. To challenge and measure the effectiveness of state-of-the-art anomaly detection for Scene AD, we introduced the ToyCity dataset. This new benchmark focuses on viewpoint variance in scenes of multiple objects and serves as a testbed for our new method OmniAD. We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset as well as the object centric established MAD-Real, establishing a new state-of-the-art for this task. Particularly, we found that the incorporation of a student attention module increases the efficacy of reverse distillation for pixel-level AD by expanding the effective receptive filed (ERF). Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD allowing for an increase in pixel level $F_{1}$ of up to 40% in many scenes. We hope that our work will catalyze further advancements, spurring the creation of advanced anomaly detection techniques capable of managing the complexities associated with varying viewpoints in 3D scenes.
8 Acknowledgments
The authors acknowledge partial financial support from the Texas Department of Transportation (Grant No. 0-7181), and Department of Defense (Project No. G0511607). The contents of this paper reflect the views of the authors, who are responsible for the facts and the accuracy of the data presented herein. The contents do not necessarily reflect the official views or policies of the sponsors. The authors acknowledge the use of the Carya Cluster and the advanced support from the Research Computing Data Core at the University of Houston to carry out the research presented here.